OpenStudy (anonymous):

How many solutions (using only nonnegative integers) are there to the following equation? x1 + x2 + x3 + x4 + x5 = 33

OpenStudy (rational):

familiar with stars and bars ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} x+x^2+x^3+x^4+x^5=33\\~\\ \left(x+x^2+x^3+x^4+x^5\right)\pmod 3=33 \pmod 3\\~\\ \left(x+x^2+x^3+x^4+x^5\right)\pmod 3=0\\~\\ \normalsize \text{for x=0}\\~\\ \left(x+x^2+x^3+x^4+x^5\right)\pmod 3=0\\~\\ 0\pmod 3\neq 0\\~\\ \normalsize \text{but }\\~\\ 0+0^2+0^3+0^4+0^5\neq 33\\~\\ \normalsize \text{for x=1}\\~\\ \left(x+x^2+x^3+x^4+x^5\right)\pmod 3=0\\~\\ 2\pmod 3\neq 0\\~\\ \normalsize \text{for x=2}\\~\\ \left(x+x^2+x^3+x^4+x^5\right)\pmod 3=0\\~\\ 2\pmod 3\neq 0\\~\\ \normalsize \text{no integer solutions}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{basic approach}\\~\\ x^1 + x^2 + x^3 + x^4 + x^5 = 33\\~\\ \normalsize \text{for x=1}\\~\\ x^1 + x^2 + x^3 + x^4 + x^5\\~\\ =5\\~\\ 5<33\\~\\ \normalsize \text{for x=2}\\~\\ x^1 + x^2 + x^3 + x^4 + x^5\\~\\ =2^1 + 2^2 + 2^3 +2^4 + 2^5\\~\\ =62\\~\\ 33<62\\~\\ \normalsize \text{again no integer solutions}\\~\\ \end{align}}\)

OpenStudy (amilapsn):

If you meant: \(x_1+x_2+x_3+x_4+x_5=33\) then there will be 66045 answers....

OpenStudy (amilapsn):

Ex: \(x_1,x_2,x_3,x_4=0,x_5=33\),etc....